yah I have one doubt 
How can i solve this problem
A galvanometer having 30 divisions has cureent sensitivity of 20 mu A/division. It has a resistance of 25 ohms.How will you convert it into an ammeter measuring upto 1A?How will you now convert this ammeter into a voltmeter reading upto 1v?

Grade:12th pass

1 Answers

Arun
4 years ago
number of divisions =30 
current senstivity = 20 microampere/division
Maximum Current flowing through galvanometer, I= 30*20microampere
Resistance, R = 25 ohm
Current required, I`=1A
to convert galvanometer into ammeter a resistor is applied in parallel  to it. and hence potential difference remains same for both i.e. galvanometer and resistor
IR= I`R`
600*10(-6) * 25= 1* R`
R` = 0.015 ohm